# SemSteDiff: Generative Diffusion Model-based Coverless Semantic Steganography Communication

### v1

@@ VI Conclusions
In this paper, we have introduced SemSteDiff to achieve coverless SemSteCom. Based on diffusion model, SemSteCom employs a semantic-key controlled conditional latent diffusion process to generate coverless stego image. Only legitimate receivers with both private and public keys enable to decode the correct secret image. Experimental results have demonstrated its effectiveness in guaranteeing semantic recovery of secret image for legitimate receivers and defending against intelligent eavesdroppers. In the future, we will further consider the integration between the conditional diffusion model and JSCC framework. By jointly optimizing both stego image generation and semantic transmission, we will focus on developing an end-to-end scheme to achieve lightweight and effective coverless semantic steganography communication.

### v2

@@ VI Conclusions
In this paper, we have introduced SemSteDiff to achieve coverless SemSteCom. Based on diffusion model, SemSteCom employs a semantic-key controlled conditional latent diffusion process to generate coverless stego image. Only legitimate receivers with both private and public keys enable to decode the correct secret image. Experimental results have demonstrated its effectiveness in guaranteeing semantic recovery of secret image for legitimate receivers and defending against intelligent eavesdroppers. In the future, we will devote to mathematical derivation of the semantic steganographic capacity, providing systematically definition and capacity boundaries. Meanwhile, we will further consider using lightweight LLMs fine-tuned specifically for key generation, or applying knowledge distillation to extract a task-specific lightweight model from a large-scale LLM, to improve the deployment consumption. Adopting advanced sampling method such as one-step-sampling is also put into consideration to reduce the generation latency. Furthermore, by jointly optimizing both stego image generation and semantic transmission, we will focus on developing an end-to-end scheme to achieve lightweight and effective coverless semantic steganography communication.